# A Collision-Free Hot-Tier Extension for Engram-Style Conditional Memory: A Controlled Study of Training Dynamics

### v1

@@ 6.2 Collisions as Implicit Regularization: Mechanisms and Boundaries
Collisions may provide an implicit regularization effect. Mechanistically, collisions cause high-frequency n-gram embeddings to be “averaged”—semantically similar but not identical n-grams share the same vector, equivalent to soft clustering. This reduces the model’s risk of overfitting to specific n-grams. Conversely, the collision-free hot tier lets each embedding serve only a single n-gram, making overfitting easier.
From an information-theoretic perspective, collisions limit the effective capacity of the embedding table: even if the table has $M$ slots, due to collisions, the actual number of distinguishable n-grams is less than $M$. This forces the model to learn more generalizable representations rather than memorizing specific n-grams’ idiosyncratic patterns. This is similar to the regularization principle of bottleneck structures (such as VAE’s latent space).
Observations supporting this explanation are: Hash configurations’ flip is approximately 250–1000 iterations later than Nine, and the cold advantage after flipping is smaller—consistent with the appearance of classic regularization techniques (such as dropout, label smoothing) delaying overfitting. However, the above observations are only correlational evidence; this paper has not directly verified the causal relationship between collisions and regularization.
However, collisions as regularization have their boundary conditions:
@@
  • Collision degree: At the two table_size endpoints we tested (300K and 800K), flip timing is identical (iter 3000), suggesting the regularization effect of collisions may have a threshold, beyond which the specific degree has little impact. But this paper only compared two endpoints; denser sweeps are needed for verification.
  • N-gram semantic distribution: The “soft clustering” effect of collisions depends on which n-grams collide. If colliding n-grams are semantically similar, averaging is beneficial; if semantically opposite, it introduces noise. Current random hashing cannot control this.
@@ 6.3 Gating Credit Assignment: Why Preferences Fixate
Our proposed “preference fixation hypothesis” can be further explained from an optimization dynamics perspective. Gating weight $\alpha$ is determined by the dot product of hidden state $h_{t}$ and key $k_{t}$, where $k_{t}=W_{K}\cdot e_{t}$ depends on memory embedding $e_{t}$. In early training, hot embeddings converge quickly due to high-frequency samples, forming stable $k_{t}$ representations; $W_{K}$ correspondingly learns to “prefer” these stable keys. Once the parameter space of $W_{K}$ is occupied by this preference, subsequent gradient updates have difficulty reversing it—because cold tier embeddings continue to change due to collisions, making it hard for $W_{K}$ to form stable preference for them.
@@ 6.4 Dialogue with the Original Engram Design
DeepSeek chose multi-head hashing rather than MPHF or other collision-free schemes in the original Engram design. The original paper explicitly acknowledges the collision problem and handles it through two mechanisms: (1) using multi-head hashing to “mitigate collisions”; (2) using context-aware gating to suppress collision-induced noise—the paper states that when retrieved memory contradicts the current context, the gate approaches zero, “effectively suppressing noise.” However, the original paper positions collisions as noise to be suppressed, not as potentially beneficial properties. Reasons for choosing multi-head hashing over collision-free schemes may include: (1) engineering simplicity—no need to pre-build static key sets; (2) dynamism—MPHF only supports static sets, while multi-head hash naturally supports arbitrary inputs.
@@ 6.5 Recommendations for Practitioners
At the scale tested in this paper (100M tokens, 125M parameters), retaining collisions is the more robust choice. Specifically: (1) when training data is relatively limited (e.g., <<1B tokens), the regularization effect of collisions helps prevent overfitting; (2) when model scale is smaller, overfitting risk is higher, and the implicit constraint from collisions is more valuable; (3) retaining collisions avoids introducing additional components like MPHF, making engineering simpler; (4) no membership test needed, lower inference latency. Whether eliminating collisions has value at larger scales or in specific domains requires future research validation (see Section 7.2).
@@ 7.1 Limitations
This paper’s experiments have the following limitations:
Scale and Architecture Differences
This paper’s experiments were conducted at 125M parameters, 100M tokens scale, while DeepSeek’s original Engram was trained on tens of billions of parameters, hundreds of billions of tokens in an MoE architecture. We use a simple GPT-2 dense architecture and cannot reproduce the interaction effects between Engram and MoE—MoE’s sparse activation may form some synergy or competition with Engram’s n-gram routing. Additionally, at larger scales, the regularization effect of collisions may weaken, and the timing and magnitude of hot/cold flipping may differ.
@@
We only test the GPT-2 architecture; other architectures like Llama have different attention and FFN designs, which may interact differently with the Engram module.
@@ 7.2 Future Work
Gating preference fixation is one of the core issues revealed in this paper. Improvement directions include: (1) enhancing gating input—letting the gate see more information, such as explicit is_hot flags, n-gram frequency, collision degree estimates, or membership confidence; (2) decoupling hot/cold gating parameters—using independent $W_{K}$ for both paths, avoiding early preference fixation affecting late adaptation; (3) introducing dynamic calibration mechanisms—such as periodic reset of gating parameters, or using EMA decay to let early preferences gradually “forget”; (4) designing auxiliary losses—explicitly penalizing mismatch between gating weights and actual prediction performance.
@@
This paper’s 100M token experiment is in a “relatively data-scarce” regime, which may amplify the regularization effect of collisions. A key question is: at 10B+ tokens, 1B+ parameter scale, does the regularization value of collisions remain significant? Additionally, different domains’ n-gram distribution differences may lead to different optimal strategies—code’s n-gram semantics are more regular, collision noise costs may be higher; while natural language’s ambiguity may make collision’s “soft clustering” more valuable. Systematic cross-scale, cross-domain experiments will help clarify the applicability boundaries of collision effects.
@@ 8 Conclusion
This paper answers a question that seems intuitively obvious but lacks empirical evidence: does eliminating collisions for high-frequency n-grams in Engram-style hashed memory improve training outcomes? The answer is no.
Under strict iso-parameter control, the collision-free Engram-Nine did not significantly outperform the collision-prone Hash baseline—although Nine-100/400K achieved the lowest validation loss, its advantage is only 0.001, far smaller than measurement standard deviation. More surprisingly, the highest-collision Hash-300K performed comparably to the collision-free Nine-100/400K. This result directly challenges the intuitive assumption that “collisions are the main bottleneck.”
Through route-stratified evaluation, we reveal the mechanism behind this phenomenon: collisions act as implicit regularization. All configurations experience a hot-to-cold performance flip—early hot positions are easier to predict, but late cold positions have lower loss. The key observation is: collision-free Nine flips earlier (iter 2000–2750), collision-prone Hash flips later (iter 3000). Collisions, by forcing semantically similar n-grams to share embeddings, play a dropout-like regularization role, delaying hot tier overfitting.
We also discovered the gating mechanism’s credit assignment problem: the gate learns to “prefer hot” in early training, but this preference fails to adjust in time after the hot/cold flip, causing high $\alpha$ positions to actually have higher loss. This mismatch suggests: the current architecture’s true bottleneck may not be in index precision, but in how gating assigns trust.
This paper’s core warning is: improving precision does not equal improving training benefits. Before pursuing “more precise” retrieval, one needs to first verify whether precision is truly the bottleneck—otherwise, added system complexity is not only unhelpful but may destroy beneficial properties implicit in the original design.

### v2

@@ 6.2 Collisions as Implicit Regularization: Mechanisms and Boundaries
Hypothesized mechanism.
Collisions may provide an implicit regularization effect through two complementary pathways. First, collisions cause high-frequency n-gram embeddings to be “averaged”—semantically similar but not identical n-grams share the same vector, equivalent to soft clustering. This reduces the model’s risk of overfitting to specific n-grams. Conversely, the collision-free hot tier lets each embedding serve only a single n-gram, making overfitting easier. Second, from an information-theoretic perspective, collisions limit the effective capacity of the embedding table: even if the table has $M$ slots, due to collisions, the actual number of distinguishable n-grams is less than $M$. This forces the model to learn more generalizable representations rather than memorizing specific n-grams’ idiosyncratic patterns—similar to the regularization principle of bottleneck structures (such as VAE’s latent space).
Supporting observations.
Hash configurations’ flip is approximately 250–1000 iterations later than Nine, and the cold advantage after flipping is smaller. This pattern is consistent with classic regularization techniques (such as dropout, label smoothing) that delay overfitting onset.
Limitations and boundary conditions.
The above observations are correlational evidence; this paper has not directly verified the causal relationship between collisions and regularization. Moreover, collisions as regularization likely have boundary conditions:
@@
  • Collision degree: At the two table_size endpoints we tested (300K and 800K), flip timing is identical (iter 3000), suggesting the regularization effect of collisions may have a threshold, beyond which the specific degree has little impact. Denser sweeps are needed for verification.
  • N-gram semantic distribution: The “soft clustering” effect depends on which n-grams collide. If colliding n-grams are semantically similar, averaging is beneficial; if semantically opposite, it introduces noise. Current random hashing cannot control this.
@@ 6.3 Gating Credit Assignment: Why Preferences Fixate
The “preference fixation hypothesis” proposed there can be further explained from an optimization dynamics perspective. Gating weight $\alpha$ is determined by the dot product of hidden state $h_{t}$ and key $k_{t}$, where $k_{t}=W_{K}\cdot e_{t}$ depends on memory embedding $e_{t}$. In early training, hot embeddings converge quickly due to high-frequency samples, forming stable $k_{t}$ representations; $W_{K}$ correspondingly learns to “prefer” these stable keys. Once the parameter space of $W_{K}$ is occupied by this preference, subsequent gradient updates have difficulty reversing it—because cold tier embeddings continue to change due to collisions, making it hard for $W_{K}$ to form stable preference for them.
@@ 6.4 Dialogue with the Original Engram Design
DeepSeek chose multi-head hashing rather than MPHF or other collision-free schemes in the original Engram design. The original paper explicitly acknowledges the collision problem and handles it through two mechanisms: (1) using multi-head hashing to “mitigate collisions”; (2) using context-aware gating to suppress collision-induced noise—the paper states that “if the retrieved memory contradicts the current context, the gate tends toward zero, effectively suppressing the noise.” However, the original paper positions collisions as noise to be suppressed, not as potentially beneficial properties. Reasons for choosing multi-head hashing over collision-free schemes may include: (1) engineering simplicity—no need to pre-build static key sets; (2) dynamism—MPHF only supports static sets, while multi-head hash naturally supports arbitrary inputs.
@@ 6.5 Recommendations for Practitioners
At the scale tested in this paper (100M tokens, 185M backbone parameters), retaining collisions is the more robust choice. Specifically: (1) when training data is relatively limited (e.g., <<1B tokens), the regularization effect of collisions helps prevent overfitting; (2) when model scale is smaller, overfitting risk is higher, and the implicit constraint from collisions is more valuable; (3) retaining collisions avoids introducing additional components like MPHF, making engineering simpler; (4) no membership test needed, lower inference latency. Whether eliminating collisions has value at larger scales or in specific domains requires future research validation (see Section 7.2).
@@ 7.1 Limitations
This work is a rapid proof-of-concept validation of a research idea, not a definitive conclusion. The limitations listed below mean that actual performance at production scale may differ significantly from what we observed. The primary contribution of this paper lies in proposing the Engram-Nine architecture, the route-stratified evaluation methodology, and the analytical framework for understanding hot/cold dynamics and gating credit assignment—not in the specific numerical outcomes. With sufficient resources and time for continued research [^1] Readers interested in funding or collaborating on follow-up research are welcome to contact the author., larger-scale experiments could reveal substantially different results; the collision-as-regularization effect we observed might diminish or even reverse when overfitting risk is inherently lower at scale.
Specifically, this paper’s experiments have the following limitations:
Scale Differences
This paper’s experiments were conducted at 185M backbone parameters, 100M tokens scale, while DeepSeek’s original Engram was trained on tens of billions of parameters, hundreds of billions of tokens. Additionally, our Engram embedding capacity (128M) represents $\sim$69% of the backbone (185M), a ratio likely much higher than in practical large-scale deployments where Engram serves as a lightweight auxiliary. This may amplify overfitting tendencies and the regularization benefit of collisions. At larger scales with more balanced Engram-to-backbone ratios, the collision effects we observed may differ significantly.
@@
We only test the GPT-2 dense architecture. DeepSeek’s original Engram uses an MoE architecture, where sparse activation may form some synergy or competition with Engram’s n-gram routing.
@@ 7.2 Future Work
Gating preference fixation is one of the core issues revealed in this paper. Improvement directions include: (1) enhancing gating input—letting the gate see more information, such as explicit hot-tier indicators, n-gram frequency, collision degree estimates, or membership confidence; (2) decoupling hot/cold gating parameters—using independent $W_{K}$ for both paths, avoiding early preference fixation affecting late adaptation; (3) introducing dynamic calibration mechanisms—such as periodic reset of gating parameters, or using EMA decay to let early preferences gradually “forget”; (4) designing auxiliary losses—explicitly penalizing mismatch between gating weights and actual prediction performance.
@@
This paper’s 100M token experiment is in a “relatively data-scarce” regime, which may amplify the regularization effect of collisions. A key question is: at 10B+ tokens, 1B+ parameter scale, does the regularization value of collisions remain significant? Relatedly, our near 1:1 Engram-to-backbone parameter ratio is likely unrealistic for production systems; exploring smaller ratios (e.g., 1:4, 1:8) may reveal different collision dynamics or even benefits from collision-free designs when overfitting risk is inherently lower. Additionally, different domains’ n-gram distribution differences may lead to different optimal strategies—code’s n-gram semantics are more regular, collision noise costs may be higher; while natural language’s ambiguity may make collision’s “soft clustering” more valuable. Systematic cross-scale, cross-domain experiments will help clarify the applicability boundaries of collision effects.
@@ 8 Conclusion
This paper investigates a question that seems intuitively obvious but lacks empirical evidence: does eliminating collisions for high-frequency n-grams in Engram-style hashed memory improve training outcomes? At the scale we tested, the answer appears to be no. Under strict iso-parameter control, collision-free indexing did not yield significant improvements over collision-prone baselines.
Through route-stratified evaluation, we uncover two key findings. First, collisions appear to act as implicit regularization: collision-free configurations exhibit earlier overfitting on high-frequency positions, while collision-prone configurations delay this effect. Second, we identify a gating credit assignment problem: the gate learns early preferences that fail to adapt as training dynamics shift, creating a mismatch between gating weights and actual prediction performance.
The central implication is: improving precision does not necessarily improve training benefits. Before pursuing “more precise” retrieval, one should first verify whether precision is truly the bottleneck—otherwise, added complexity may destroy beneficial properties implicit in the original design. We hope the route-stratified evaluation methodology and the analytical framework proposed here will prove useful for diagnosing similar systems.